Find the power of i.
-17
i

Answer by khwang(438)

Be careful about your notation. It seems that you should type as
i^(-17)
Using exponential law,

i^(-17) = (i^17)^(-1) = (i^16 * i)^(-1) [since i^16= (i^4)^4 & i^4 = 1 ]
= i^(-1) = 1/i = -i. [since i^2 = -1]
Kenny